OpenStudy (anonymous):

When the quadratic equation y = (x − 3)(4x + 2) is written in standard form, what is the value of the coefficient "b?"

OpenStudy (anonymous):

well, have u tried it yet :) ?

OpenStudy (anonymous):

yea i think its −6

OpenStudy (anonymous):

can u show ur work? i think there is a little mistake somewhere...thats the value of coefficient c

OpenStudy (anonymous):

oh lets me try again

OpenStudy (anonymous):

good, ok :)

OpenStudy (anonymous):

y = (x − 3)(4x + 2) y = 4x² + 2x − 12x − 6 y = 4x² − 10x − 6 y = ax² + bx + c ⇒ b = −10

OpenStudy (anonymous):

so its -10

OpenStudy (anonymous):

quite right :)

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

np :)